# Near-Optimal Emission-Aware Online Ride Assignment Algorithm for Peak Demand Hours

### v1

@@ 5 Experimental Analysis
In this section, we evaluate the performance of LARA using both synthetic data and real-world data from the RideAustin dataset. This dual approach allows us to compare LARA against alternative algorithms under controlled conditions as well as real-world ridesharing scenarios. Synthetic datasets help us to better simulate the high demand hours and analyze the impact of various factors—such as average trip distances, the number of drivers, and the time interval between consecutive batch assignments—on the performance of LARA and the baseline algorithms. Additionally, the evaluation on the RideAustin dataset provides insights into how LARA performs in real-world conditions. In all experimental analyses, we use the batch system described in Section 2.1 to perform the ride assignments.
@@ 5.1 Experiments on Synthetic Data
Comparison Algorithms. We compare the performance of LARA against two algorithms: (1) TORA [30], the only existing emission-aware ride assignment algorithm; and (2) a heuristic algorithm that assigns riders to the closest available driver, referred to as CD.
Experimental Setup. We generated traces of 10,000 ride requests, where a new ride request is posted on the ridesharing platform every 30 seconds. Pickup locations are randomly selected within Austin, Texas, and drop-off locations are also randomly chosen such that the average trip distance is 15 km, with a standard deviation of 5 km. The ridesharing platform operates with 120 drivers, whose vehicle emissions range between 70–300 gCO2/km. This range mirrors the emissions of vehicles found in the RideAustin dataset and has been used in previous studies [29, 30].
For LARA, we used deadhead distance limits options of 1, 2, 5, 10, 15, and 30 km, with $\alpha=0.75$ and $Q_{\max}=40$, which encourages LARA to limit the portion of waiting time spent in the assignment queue to under 20 minutes. This is achieved as four new ride requests are generated in each batch, and 10 batch assignments occur within a 20-minute window. Additionally, we applied the Monte Carlo estimation method to estimate the function $g_{M}$. When LARA assigns a ride while limiting the deadhead distance to $d$, it calculates the emission reduction compared to the closest driver’s emission and updates the estimate of $g_{M}(d)$ accordingly. In each evaluation, all parameters are kept constant except for one, which is varied to assess its impact. For each test, we report the average trip emissions and rider waiting times for LARA and the comparison algorithms.
Experiment Results. Figure LABEL:fig:synthetic_batchDur illustrates the average trip emissions and rider waiting times as a function of batch duration for LARA and the comparison algorithms. As seen in the figure, LARA consistently achieved the lowest average emissions and waiting times across all batch durations. Specifically, compared to CD, LARA reduced the average trip emissions by $30.1\%-34.2\%$, while this range for TORA was $24.1\%-26.2\%$. In terms of waiting times, LARA resulted in average waiting times between $1532$ and $1640$ seconds, whereas for CD, the range was $1665$ to $2150$ seconds, and for TORA, between $1931$ and $3164$ seconds. During high-demand hours, the number of available drivers is low, leaving algorithms like CD and TORA with less flexibility to reduce emissions, as they tend to assign riders to any available driver. In contrast, LARA prioritizes finding rides with lower emissions and shorter deadhead distances, leading to lower average emissions and waiting times overall.
@@
In the second evaluation, we examine how the number of drivers in the ridesharing platform affects the performance of online ride assignment algorithms. Figure LABEL:fig:synthetic_driver illustrates the average trip emissions and rider waiting times as a function of the number of drivers for LARA and comparison algorithms. Our results show a significant performance gap between LARA and the comparison algorithms when the number of drivers is low. In such scenarios, demand exceeds driver availability, which amplifies the performance difference between the optimal algorithm and alternatives like TORA and CD. For instance, with 80 drivers, LARA achieves $31\%$ lower emissions compared to TORA and reduces waiting times by $42.9\%$. On the other hand, as the number of drivers increases, the performance gap between the algorithms narrows. With larger driver pools (e.g., more than 125 drivers in our tests), the platform is no longer in a high-demand state, allowing algorithms like CD to achieve the shortest waiting times. Nonetheless, even in these conditions, LARA continues to deliver the lowest average emissions.
@@ 5.2 Experiments on Real Dataset
In this section, we evaluate the performance of LARA and the comparison algorithms using a real-world dataset. We utilized ride request data from the RideAustin dataset [25] for the first week of December 2016, which contains $29850$ ride requests. To simulate high-demand conditions, we limited the number of available drivers to the first 50 from the dataset. As in the synthetic experiments, we tested deadhead distance limits of 1, 2, 5, 10, 15, and 30 km, set $Q_{\max}=40$, and evaluated three versions of LARA with $\alpha$ values of 0.25, 0.5, and 0.75.
We report the average emissions per trip and the average rider waiting times for various batch durations. The results are shown in Figure 4. As illustrated, increasing the batch duration leads to lower average emissions but longer waiting times. This trend is consistent with the results from the synthetic dataset (Figure LABEL:fig:synthetic_batchDur). Across all batch durations, the three versions of LARA outperformed the comparison algorithms, achieving lower average emissions and shorter waiting times. Specifically, compared to CD, LARA with $\alpha=0.75$, $\alpha=0.5$, and $\alpha=0.25$ reduced emissions up to $11.1\%$, $11.3\%$, and $13.9\%$, respectively, while TORA reduced emissions up to $6.6\%$. Additionally, LARA with $\alpha=0.25$ achieved between $8.7\%$ and $15.7\%$ shorter waiting times compared to CD across the different batch durations.
Finally, we report the percentage of ride requests performed by drivers with relatively low, and high emission vehicles to assess the performance of LARA and other comparison algorithms in providing fair ride assignment across different vehicles. To this end, we categorize vehicles with unit emission of less than $2150gCO$ 2 as a low emission vehicles and vehicles with unit emission of higher than $2250gCO$ 2 a high emission vehicles. Using this categorization approach, $24\%$ of dataset drivers categorized to low emission vehicles and $26\%$ categorized to high emission vehicles. The fraction of ride requests performed by low/high emission vehicles must be very similar to the fraction of those vehicles in the platform under control of fully fair algorithm.
In Table 1, and Table 2 we present the percentage of ride requests fulfilled by drivers of low- and high-emission vehicles, along with the percentage of their total distance spent as deadhead miles. The results indicate that algorithms like LARA with high $\alpha$ values, or TORA, assign fewer rides to low-emission vehicles. This occurs because these algorithms prioritize minimizing deadhead distances over reducing emissions, causing low-emission vehicles to spend more of their total distance on deadhead miles. For example, under ride assignment of TORA, $55.4\%$ of the distance traveled by low-emission vehicles was due to deadhead miles, compared to $49.2\%$ under LARA with $\alpha=0.25$ and $51.8\%$ under CD. During high-demand periods, when all drivers are consistently busy, low-emission vehicles tend to travel more deadhead miles, resulting in fewer ride assignments. This underscores the trade-off between fair ride allocation and optimizing carbon reduction in ridesharing platforms.
@@ 6 Related Work
Recent advancements in ridesharing services have led to significant developments through various approaches. Theoretical approaches including the creation of formal problem models and the development of advanced solution algorithms. Empirical approaches provide valuable insights, while machine learning techniques enhance these insights with sophisticated, data-driven solutions, together enabling more effective and refined advancements in the field.
On the theoretical side, significant effort is invested in modeling and understanding the structure of the ridesharing process, as this is essential for improving platform efficiency. Sadowsky et al. [28] investigate the impact of Uber and Lyft’s entry into major US urban areas on public transportation usage. The study uses a regression discontinuity design to assess how these ridesharing services influence public transit ridership. For example, Afeche et al. [1] develop and analyzed an integrated, game-theoretic model that characterizes the system equilibrium in ridesharing platforms. Hu et al. [16] study surge pricing strategies in ridesharing platforms by controlling rider fares and driver wages to balance demand with supply. The main goal of the work by Castillo et al.  [6] is to demonstrate how surge pricing can improve efficiency in ridesharing platforms by reducing the likelihood of an inefficient ”wild goose chase.” Their stylized model highlights how surge pricing reduces long pickup times, which in turn helps increase driver earnings by optimizing the allocation of idle drivers. Garg et al. [15] propose driver payment mechanisms for ridesharing platforms, focusing on enhancing the incentive compatibility of surge pricing and studying such mechanisms under non-stationary demand conditions. Ma et al. [21] propose a spatio-temporal pricing mechanism designed to maximize welfare by aligning incentives. This mechanism sets dynamic prices that are smooth both spatially and temporally, ensuring that drivers accept their assigned trips instead of seeking alternative opportunities. Feng et al. [12] investigate a two-stage stochastic matching model for ridesharing platforms, focusing on the challenges posed by uncertainty in rider and driver availability. Bistaffa et al. [3] adopt a cooperative game theoretic approach in order to tackle the social ridesharing problem by optimizing coalition formation to reduce overall travel costs. Vazifeh et al. [33] propose an optimization-based method by developing a network-based solution to calculate the minimal number of vehicles required to fulfill a set of trips without causing delays. Bian et al. [2] investigate how to optimally match vehicles with passengers based on individual personalized requirements in ridesharing platforms.
Research on the empirical and machine learning aspects of ridesharing focuses on utilizing data-driven methods and machine learning to reveal patterns in the ridesharing process to improve various aspects of ridesharing services. Chen et al. [9]demonstrate how surge pricing encourages drivers to extend their working hours, thereby enhancing overall efficiency. Bongiovanni et al. [5] presents a machine learning-based optimization framework for autonomous ridesharing operations, using a two-phase approach to improve vehicle-trip assignments. Fielbaum et al. [13]investigate the sources of unreliability in ridesharing systems, focusing on how user-to-vehicle assignments impact trip delays and waiting times, using a New York City case study to assess trade-offs between uncertainty reduction and service quality. Riley et al. [26] study optimizing large-scale real-time ridesharing systems by reducing customer wait times and minimizing the number of vehicles required, while maintaining service guarantees and adhering to ride duration limits. Their approach integrates a dispatching algorithm, a machine learning model to predict demand, and model predictive control to reposition idle vehicles. Yatnalkar et al. [36] present an Enhanced ridesharing Model (ERSM) that leverages machine learning to match riders based on human characteristics and predict their key preferences, with the goal of enhancing user satisfaction. The model tackles social barriers, long wait times, and pricing challenges in ridesharing. Tong et al. [32] propose a unified linear regression model with 200 million multidimensional features, along with optimization strategies to enhance the efficiency of model training and updates. Since ridesharing demand depends heavily on supply, the observed demand in historical data reflects only a portion of the true demand, limited by available supply. The primary goal of their work is to create a multi-community spatio-temporal graph convolutional network framework for accurately predicting passenger demand across multiple regions in urban transportation systems. Fielbaum et al. [13]investigate the sources of unreliability in ridesharing systems, focusing on how user-to-vehicle assignments impact trip delays and waiting times, using a New York City case study to assess trade-offs between uncertainty reduction and service quality. Riley et al. [26] study optimizing large-scale real-time ridesharing systems by reducing customer wait times and minimizing the number of vehicles required, while maintaining service guarantees and adhering to ride duration limits. Their approach integrates a dispatching algorithm, a machine learning model to predict demand, and model predictive control to reposition idle vehicles. Yatnalkar et al. [36] present an Enhanced ridesharing Model (ERSM) that leverages machine learning to match riders based on human characteristics and predict their key preferences, with the goal of enhancing user satisfaction. The model tackles social barriers, long wait times, and pricing challenges in ridesharing. Tong et al. [32] propose a unified linear regression model with 200 million multidimensional features, along with optimization strategies to enhance the efficiency of model training and updates. Since ridesharing demand depends heavily on supply, the observed demand in historical data reflects only a portion of the true demand, limited by available supply. The primary goal of their work is to create a multi-community spatio-temporal graph convolutional network framework for accurately predicting passenger demand across multiple regions in urban transportation systems. Liu et al. [20] develop a personalized ridesharing demand prediction model that learns unique weights to capture distinct patterns and variations in demand across various locations and times.
@@ 8 Acknowledgment
Appendix A Proof of Theorem 1
Proof. We begin by defining a Lyapunov function $L(b)$ and the conditional Lyapunov drift function, $D(b)$, as follows:
$L(b)=$ $\frac{1}{2}Q(b)^{2},$
$D(b)=$ $\mathbb{E}\left[L(b+1)-L(b)|Q(b),M_{b}\right].$
Next, we evaluate the drift function using the update rule for $Q(b)$ from Equation (6).
$Q(b+1)^{2}=$ $Q(b)^{2}+r_{b}^{2}+\sum_{d\in\mathcal{D}}a_{b,d}\cdot\beta_{b}^{2}-2r_{b}Q(b)$
$+$ $2Q(b)\sum_{d\in\mathcal{D}}a_{b,d}\cdot\beta_{b}-2r_{b}\sum_{d\in\mathcal{D}}a_{b,d}\cdot\beta_{b}.$
$\Rightarrow D(b)=$ $\frac{1}{2}\mathbb{E}[(r_{b}-\sum_{d\in\mathcal{D}}a_{b,d}\cdot\beta_{b})^{2}|Q(b),M_{b}]$
-- $Q(b)\mathbb{E}[(r_{b}-\sum_{d\in\mathcal{D}}a_{b,d}\cdot\beta_{b})|Q(b),M_{b}].$
By subtracting a value of $Q_{\max}\cdot\beta_{b}\sum_{d\in\mathcal{D}}a_{b,d}\cdot(1-\alpha+\alpha\cdot g_{M}(d))$ from both sides, we obtain:
$D(b)-$ $Q_{\max}\cdot\beta_{b}\sum_{d\in\mathcal{D}}a_{b,d}\cdot(1-\alpha+\alpha\cdot g_{M}(d))$
$=$ $\frac{1}{2}\mathbb{E}[(r_{b}-\sum_{d\in\mathcal{D}}a_{b,d}\cdot\beta_{b})^{2}|Q(b),M_{b}]$
-- $Q(b)\mathbb{E}[(r_{b}-\sum_{d\in\mathcal{D}}a_{b,d}\cdot\beta_{b})|Q(b),M_{b}]$
-- $Q_{\max}\cdot\beta_{b}\sum_{d\in\mathcal{D}}a_{b,d}\cdot(1-\alpha+\alpha\cdot g_{M}(d))$
$\leq$ $\frac{1}{2}\mathbb{E}[(r_{b}-\sum_{d\in\mathcal{D}}a_{b,d}\cdot\beta_{b})^{2}|Q(b),M_{b}]$
-- $Q(b)\mathbb{E}[(r_{b}-\psi\sum_{d\in\mathcal{D}}a^{*}_{b,d}\cdot\beta_{b})|Q(b),M_{b}]$
-- $\psi Q_{\max}\cdot\beta_{b}\sum_{d\in\mathcal{D}}a^{*}_{b,d}\cdot(1-\alpha+\alpha\cdot g_{M}(d))$
$\leq$ $\frac{1}{2}\mathbb{E}[(r_{b}-\sum_{d\in\mathcal{D}}a_{b,d}\cdot\beta_{b})^{2}|Q(b),M_{b}]$
$+$ $Q(b)\mathbb{E}[\sum_{d}a_{b,d}(d+d_{t,b})]\mathbb{E}[(\sum_{d\in\mathcal{D}}a^{*}_{b,d}\frac{\beta_{b}}{(d+d_{t,b})}$
-- $\frac{r_{b}}{\sum_{d}a_{b,d}(d+d_{t,b})})|Q(b),M_{b}]$
-- $\psi Q_{\max}\cdot\beta_{b}\sum_{d\in\mathcal{D}}a^{*}_{b,d}\cdot(1-\alpha+\alpha\cdot g_{M}(d)),$
The second term is always negative, as the expected rate of assignment under the optimal stationary algorithm cannot exceed the expected posting rate of new ride requests. By taking the conditional expectation of both sides, summing over all batches, and dividing by $Q_{\max}\cdot B_{N}\mathbb{E}[d+d_{t,b}]$, we get:
$\Rightarrow$ $\frac{1}{Q_{\max}\cdot B_{N}\mathbb{E}[d+d_{t,b}]}\sum_{b=1}^{B_{N}}D(b)-\sum_{b=1}^{B_{N}}\beta_{b}\sum_{d\in\mathcal{D}}a_{b,d}\frac{1-\alpha+\alpha\cdot g_{M}(d)}{B_{N}\mathbb{E}[d+d_{t,b}]}$
$\leq$ $\frac{1}{2Q_{\max}\cdot B_{N}\mathbb{E}[d+d_{t,b}]}\sum_{b=1}^{B_{N}}\mathbb{E}[(r_{b}-\sum_{d\in\mathcal{D}}a_{b,d}\cdot\beta_{b})^{2}|Q(b),M_{b}]$
-- $\sum_{b=1}^{B_{N}}\beta_{b}\sum_{d\in\mathcal{D}}a^{*}_{b,d}\frac{1-\alpha+\alpha\cdot g_{M}(d)}{B_{N}\mathbb{E}[d+d_{t,b}]}$
$\leq$ $\frac{\mathbb{E}[r_{b}^{2}]}{Q_{\max}\cdot\mathbb{E}[d+d_{t,b}]}-\sum_{b=1}^{B_{N}}\beta_{b}\sum_{d\in\mathcal{D}}a^{*}_{b,d}\frac{1-\alpha+\alpha\cdot g_{M}(d)}{B_{N}\mathbb{E}[d+d_{t,b}]}.$
Taking the limit as $B_{N}\rightarrow\infty$ completes the proof.

### v2

@@ 5 Experimental Analysis
We begin this section with a high-level analysis of LARA’s behavior. Specifically, we examine how the parameters $Q_{\max}$ and $\alpha$ influence its decisions. We then conduct an extensive evaluation of LARA using both synthetic data and real-world data from the RideAustin dataset [28]. This dual approach allows us to compare LARA against alternative algorithms in both controlled experimental settings and realistic ridesharing environments. Synthetic datasets enable the simulation of high-demand scenarios and facilitate analysis of key factors such as trip distances, driver availability, and batch assignment intervals. Although publicly available ridesharing datasets are limited and often lack critical information (e.g., pickup/drop-off locations or vehicle attributes), the RideAustin dataset remains the most suitable for evaluating LARA. While it does not fully reflect high-demand conditions, it provides valuable insights into the algorithm’s real-world performance. Finally, our ride-assignment process follows a batch system, where the algorithm assigns unallocated requests to available drivers in successive batches. Unassigned requests are carried over to the next cycle for reassignment.
5.1 Understanding of the Behavior of LARA
First, we aim to scrutinize the deadhead control process within LARA. LARA dynamically selects the deadhead distance limit based on factors such as the number of requests in the assignment queue and hyper-parameters like $Q_{\max}$. To analyze the impact of these factors on LARA’s decision-making, we conduct a simple analysis using a version of LARA with $Q_{\max}=100$ and deadhead distance limit options of 1, 2, 5, 10, 15, and 30 km. We assume there are 15 ride requests in the assignment queue and evaluate how LARA selects deadhead limits by varying the number of requests and the value of $Q_{\max}$ in two separate analyses. In this analysis, we use a logarithmic form for the function $g_{M}(d)$, where $g_{M}(d)=C_{0}\cdot\log(d)$, with $C_{0}=1/\log(d_{\max})\approx 0.20$ as the normalization constant.
Figure 2(a) illustrates the influence of the number of requests in the assignment queue on the deadhead distance limit selected by LARA for different values of $\alpha$. When the queue has relatively few ride requests, LARA selects higher deadhead distance limits to prioritize reducing emissions. However, as the number of requests increases, LARA shifts focus towards reducing rider waiting times by selecting lower deadhead limits. Additionally, increasing the parameter $\alpha$ causes LARA to emphasize more on the emission term in the objective function of Equation (3a), leading to the selection of higher deadhead distance limits. In Figure 2(b), we plot the deadhead limits selected by LARA as a function of $Q_{\max}$ for different values of $\alpha$. When the number of requests in the assignment queue is fixed (15 in this analysis), increasing $Q_{\max}$ widens the gap between the assignment queue length and $Q_{\max}$. This results in LARA selecting higher deadhead distance limits as the system allows for a larger buffer to focus on emission reduction.
Key takeaway. As the number of unassigned ride requests increases, LARA lowers the deadhead distance limits to prioritize faster ride assignments. Higher values of $Q_{\max}$ leads LARA to select larger deadhead distance limits, providing greater flexibility to reduce emissions.
Choosing parameters $\alpha$ and $Q_{\max}$. In practice, the parameter $\alpha$ is chosen by stakeholders and decision makers based on their relative priorities between reducing emissions and minimizing waiting times. As shown in Figure 2, varying $\alpha$ influences the deadhead distance limits selected by LARA. The parameter $Q_{\max}$, on the other hand, reflects the platform’s tolerance for maximum allowable waiting times. Specifically, doubling $Q_{\max}$ can result in a doubling of the worst-case waiting time under adverse conditions. Selecting an appropriate value for $Q_{\max}$ should take into account the operational and statistical characteristics of the ridesharing platform, such as demand patterns, and traffic conditions.
@@ 5.2 Experiments on Synthetic Data
Comparison Algorithms. We compare the performance of LARA against two algorithms: (1) TORA [33], the only existing emission-aware ride assignment algorithm which also considers deadhead emissions; and (2) a heuristic algorithm that assigns riders to the closest available driver, referred to as CD. TORA is considered as a baseline algorithm for emission reduction while CD is a baseline for minimizing passengers’ waiting times in a greedy manner.
Experimental Setup. We generated traces consisting of 50,000 ride requests, with a new request posted on the ridesharing platform every 5 seconds. Pickup and drop-off locations are selected within area of Austin, Texas, ensuring an average trip distance of 15 km with a standard deviation of 5 km. The assignment happens every 2 minutes and the platform operates with 500 drivers, resulting in a per hour driver-to-request ratio of approximately $60\%$, aligning with observed Uber/Lyft statistics [26]. Additionally, vehicle emissions range between 70–300 gCO2/km, reflecting emission levels found in the RideAustin dataset and consistent with previous studies [32, 33].
For LARA, we used deadhead distance limits options of 1, 2, 5, 10, 15, and 30 km, with $\alpha=0.75$ and $Q_{\max}=240$, which encourages LARA to keep the portion of waiting time spent in the assignment queue to under 20 minutes. This is achieved as 24 new ride requests are generated in each batch, and 10 batch assignments occur within a 20-minute window. Additionally, we applied the Monte Carlo estimation method to estimate the function $g_{M}$. Specifically, when LARA assigns a ride while limiting the deadhead distance to $d$, it calculates the emission reduction compared to the closest driver’s emission and updates the estimate of $g_{M}(d)$ based on Monte Carlo estimation strategy. In each evaluation, all parameters are kept constant except for one, which is varied to assess its impact. For each test, we report the average trip emissions and rider waiting times for LARA and the comparison algorithms.
Experiment Results. Figure LABEL:fig:synthetic_batchDur illustrates the average trip emissions and rider waiting times as a function of batch duration for LARA and the comparison algorithms. As seen in the figure, LARA consistently achieved the lowest average emissions and waiting times across all batch durations. Specifically, compared to CD, LARA reduced the average trip emissions by $30\%-34\%$, while this range for TORA was $24.1\%-26.4\%$. In terms of waiting times, LARA resulted in average waiting times between $1500$ and $1600$ seconds, whereas for CD, the range was $1600$ to $2200$ seconds, and for TORA, between $1900$ and $3200$ seconds. During high-demand hours, the number of available drivers is low, leaving algorithms like CD and TORA with less flexibility to reduce emissions, as they tend to assign riders to any available driver. In contrast, LARA prioritizes finding rides with lower emissions and shorter deadhead distances, leading to lower average emissions and waiting times overall.
@@
In the second evaluation, we examine how the number of drivers in the ridesharing platform affects the performance of online ride assignment algorithms. Figure LABEL:fig:synthetic_driver illustrates the average trip emissions and rider waiting times as a function of the number of drivers for LARA and comparison algorithms. Our results show a significant performance gap between LARA and the comparison algorithms when the number of drivers is low. In such scenarios, demand exceeds driver availability, which amplifies the performance difference between the optimal algorithm and alternatives like TORA and CD. For instance, with 310 drivers, LARA achieves $31\%$ lower emissions compared to TORA and reduces waiting times by $43\%$. On the other hand, as the number of drivers increases, the performance gap between the algorithms narrows. With larger driver pools (e.g., more than 480 drivers in our tests), the platform is no longer in a high-demand state, allowing algorithms like CD to achieve the shortest waiting times. Nonetheless, even in these conditions, LARA continues to deliver the lowest average emissions.
@@
Overall takeaway. Various factors, such as the number of drivers, ride request rates, batch assignment frequency, and trip distances, can lead to high-demand conditions. Under such conditions, LARA outperforms CD and TORA in reducing average emissions and rider waiting times.
@@ 5.3 Experiments on Real Dataset
In this section, we evaluate the performance of LARA and baseline algorithms using the RideAustin dataset [28]. This dataset enables realistic simulations of ridesharing scenarios, allowing us to assess both average waiting times and emissions under various conditions. We use ride request data from the first week of December 2016, which includes 29,850 ride requests. To simulate high-demand conditions, we limit the number of available drivers so that the driver-to-request ratio is approximately 40%—compared to the typical 50% of real world conditions [26]. As in the synthetic experiments, we test deadhead distance limits of 1, 2, 5, 10, 15, and 30 km, set $Q_{\max}=240$, and evaluate three versions of LARA with $\alpha$ values of 0.25, 0.5, and 0.75.
We report the average emissions per trip and the average rider waiting times for various batch durations. The results are shown in Figure 4. As illustrated, increasing the batch duration leads to lower average emissions but longer waiting times. This trend is consistent with the results from the synthetic dataset (Figure LABEL:fig:synthetic_batchDur). Across all batch durations, the three versions of LARA outperformed the comparison algorithms, achieving lower average emissions and shorter waiting times. Specifically, compared to CD, LARA with $\alpha=0.75$, $\alpha=0.5$, and $\alpha=0.25$ reduced emissions up to $1.1\%, 11$$1.3\%$, and $13.9\%$, respectively, while TORA reduced emissions up to $6.6\%$. Additionally, LARA with $\alpha=0.25$ achieved between $8.7\%$ and $15.7\%$ shorter waiting times compared to CD across the different batch durations.
@@
Finally, we report the percentage of ride requests performed by drivers with relatively low, and high emission vehicles to assess the performance of LARA and other comparison algorithms in providing fair ride assignment across different vehicles. To this end, we categorize vehicles with unit emission of less than $150gCO2$ as a low emission vehicles and vehicles with unit emission of higher than $250gCO2$ a high emission vehicles. Using this categorization approach, $24\%$ of dataset drivers categorized to low emission vehicles and $26\%$ categorized to high emission vehicles. The fraction of ride requests performed by low/high emission vehicles must be very similar to the fraction of those vehicles in the platform under control of fully fair algorithm.
In Table 1, and Table 2 we present the percentage of ride requests fulfilled by drivers of low- and high-emission vehicles, along with the percentage of their total distance spent as deadhead miles. The results indicate that algorithms like LARA with high $\alpha$ values, or TORA, assign fewer rides to low-emission vehicles. This occurs because these algorithms prioritize minimizing deadhead distances over reducing total emissions of trips, causing low-emission vehicles to spend more of their total distance on deadhead miles. For example, under ride assignment of TORA, $55.4\%$ of the distance traveled by low-emission vehicles was due to deadhead miles, compared to $49.2\%$ under LARA with $\alpha=0.25$ and $51.8\%$ under CD. During high-demand periods, when all drivers are consistently busy, low-emission vehicles tend to travel more deadhead miles, resulting in fewer ride assignments. This underscores the trade-off between fair ride allocation and optimizing carbon reduction in ridesharing platforms.
@@ 6 Related Work
Recent advances in ridesharing systems have been fueled by a combination of theoretical modeling, empirical analysis, and data-driven techniques. From a theoretical standpoint, prior work has focused on modeling platform dynamics and designing mechanisms to improve operational efficiency. For example, Sadowsky and Nelson [31] study the causal effect of ridesharing platforms on public transit ridership using a regression discontinuity approach. Afeche et al. [1] develop a game-theoretic model to characterize equilibrium behavior in ridesharing platforms, while Hu et al. [13] and Castillo et al. [5] investigate surge pricing mechanisms that dynamically adjust fares and wages to balance supply and demand, thereby reducing inefficient driver movement. Other efforts propose incentive-compatible payment schemes under non-stationary conditions [11], and spatio-temporal pricing strategies that promote spatial and temporal continuity in driver engagement [23].
Environmental sustainability in ridesharing has also garnered significant attention. Several empirical studies quantify the potential emission reduction enabled by shared mobility. Jalali et al.[14] show that ridesharing platforms can reduce transportation emissions by up to 24% through increased vehicle utilization. A complementary analysis using historical data from Tokyo reveals that up to 27% of vehicle kilometers traveled could be saved through shared rides, resulting in emissions reductions of up to 84%. More recent work has focused on algorithmic approaches to minimize both in-trip and deadhead emissions[19, 34, 32, 33], while jointly optimizing service-level objectives such as rider waiting time and equitable assignment across vehicles with heterogeneous emissions profiles. These studies highlight critical trade-offs between efficiency, fairness, and sustainability.
A parallel line of work addresses matching and vehicle allocation under uncertainty. Feng et al. [9] propose a two-stage stochastic model to manage uncertain rider and driver availability. Cooperative game-theoretic approaches have been used for coalition formation [3], and network-based models aim to minimize fleet size while preserving service quality [37]. Preference-aware strategies further enhance match quality in dynamic systems [2].
Recent research also leverages machine learning to improve operational efficiency. Chen and Sheldon [7] show surge pricing boosts driver supply, and Bongiovanni et al. [4] develop a learning-based pipeline for autonomous ridesharing. Other studies address delay mitigation [10], real-time dispatching and repositioning [29], and rider satisfaction through behavior-aware modeling [41]. Spatio-temporal forecasting and personalized behavior encodings further improve demand prediction [36, 21].